OpenStudy (anonymous):

When a race begins, a racecar accelerates from zero to 200 km/h in 4.80 seconds. What is its average acceleration?

OpenStudy (anonymous):

Shouldn't This Be A Math Question?

OpenStudy (anonymous):

roughly 12 m/s²

OpenStudy (michele_laino):

by definition, the average acceleration A is: \[\Large A = \frac{{\Delta v}}{{\Delta t}} = \frac{{200 \cdot 3.6}}{{4.8}} = ...\frac{{meters}}{{{{\sec }^2}}}\] please 200 Km/hour correspond to 200*3.6 m/seconds

OpenStudy (kropot72):

200 km/h corresponds to \[\large \frac{200}{3.6}\ m/s\]

OpenStudy (michele_laino):

oops.. you are right! @kropot72 thanks! for your comment :) I'm very sorry for my error @Joshdaddy here is the right computation: \[\Large\frac{{\Delta v}}{{\Delta t}} = \frac{{200/3.6}}{{4.8}} = ...\frac{{meters}}{{{{\sec }^2}}}\]

OpenStudy (kropot72):

@Michele_Laino You're welcome :)

OpenStudy (michele_laino):

thanks! :)